OpenStudy (ivancsc1996):

Which volume would have at standard conditions 25.2 mL of gas hydrogen at 21 degrees celcious and at a presure of 1.96atm? Please help me, I got an answer of 46.02mL but my book says it is 82.3mL. According to my book standard conditions are 0 degrees celcious and 1 atm. The book is "Fundación de Química" Ralph A. Burns.

OpenStudy (ivancsc1996):

This was my procedure: P1=1.96atm; T1= 294K; V1=25.2mL; P2=1atm; T2=273K; V2=?\[\frac{ P _{1}V _{1} }{ T _{1} }=\frac{ P _{2}V _{2} }{ T _{2} } \rightarrow V _{2}=\frac{ P _{1}V _{1}T _{2} }{ T _{1}P _{2} }=\frac{ 1.96atm22.5mL273K }{ 294K1atm }=46.02mL\]

OpenStudy (anonymous):

my answer is coming out 2 b 55.29 ml

OpenStudy (anonymous):

sorry it is wrong wait

OpenStudy (anonymous):

Hey i checked my calculation again. My answer is also 46 ml

OpenStudy (anonymous):

i think the answer in the book is wrong

OpenStudy (anonymous):

wt do u think????

OpenStudy (ivancsc1996):

I think that as well, maybe if I do some research on the web I'll find there may be some issue with this edition.

OpenStudy (anonymous):

yup u can try

OpenStudy (anonymous):

the answer is 46.02

OpenStudy (ivancsc1996):

No, I am pretty sure the book is wrong, thank you very much!

OpenStudy (ivancsc1996):

Because I don't know what nR means. The way you are trying to do it which answer gives?

OpenStudy (ivancsc1996):

And I do have two systems, when the pressure is 1.96atm etc.... and when I put it back on standard conditions.

OpenStudy (anonymous):

...and also when u say standard conditions if u r refering 2 standard temperature and pressure it has to be 0k not 0 degees C....absolute zero....plz correct me if i am wrong

OpenStudy (ivancsc1996):

Oh no, in my book the standard temperature is 0 celcious.